Given initially equimolar solutions of each of the following weak acids, which solution will have the highest hydrogen ion concentration once equilibrium is established?

a. H2O

b. HClO2

c. HF

d. HClO
Answer #1

The concentrations of all given species are same.

We know that, as K a value increases dissociation of acid increases which in turn increases concentration of hydrogen ion. Hence, acid solution with highest K a value of acid will contain highest concentration of hydrogen ion.

Solution Ka
Water 1.01 \times 10 -14
HClO2 1.10 \times  10 -02
HF 6.8 \times 10 -04
HClO 3.0 \times 10 -08

From above table , it is clear that dissociation constant of HClO2 is highest and hence its solution will contain maximum concentration of hydrogen ion at equilibrium.

ANSWER : HClO2